OpenStudy (anonymous):

The parabola defined by y^(2)=4x+6 opens: a) down b) up c) to the left d) to the right

OpenStudy (anonymous):

Can somebody help me with my problem?

OpenStudy (anonymous):

Actually the standard formula of a parabola is y^2=4ax....But in your question there is an extra constant term.If your question was y^2=4x,then the parabola would open to the right since a=1.So,I think there is an error in your question.Please Check! Rate if u like tis answer.:)

OpenStudy (anonymous):

y=+x^2 opens up ie in the positive y direction so x=y^2 opens in the positive x direction ie to the right.

OpenStudy (anonymous):

so is it up or right?